OpenStudy (anonymous):

If f(x) = x(2x – 1) and g(x) = 2x – 1, find f(x) divided by g(x)

OpenStudy (anonymous):

i need answers!!!!

OpenStudy (anonymous):

Want to see your try first

OpenStudy (anonymous):

if f(x) x/3 and g(x) = 1/x find f(x) - g(x)

OpenStudy (anonymous):

Tom starts a lawnmowing business. Before he can begin, he purchases a new lawnmower. Each weekend he mows the same number of lawns and earns the same amount of money. Write an equation to show how much money he's earned at any week. x y Week Profit in Dollars 1 –90 2 –30 3 30 4 90 A. y = 60x – 90 B. y = 60x – 150 C. y = 90x + 150 D. y = 60x x's are 1,2,3,4, and the y is the bigger numbers!

OpenStudy (anonymous):

If f(x) = 4x + 3 and g(x) = 3x, evaluate f(x) multipled by g(x) for x= negative 1 over 2

OpenStudy (vishweshshrimali5):

@brittney989 , we all would be happy to help you but first please post your approach to the question. We are not allowed to give the answers directly.

OpenStudy (anonymous):

i just need help with all of them.

OpenStudy (anonymous):

This is not our work its your work

OpenStudy (anonymous):

nvm if you are wrong , tell what you did

OpenStudy (vishweshshrimali5):

Even if you were able to get one step only. Then post that step.

OpenStudy (anonymous):

There is just one step for the 1st one

OpenStudy (anonymous):

i thought it was combining like terms or multiplying

OpenStudy (vishweshshrimali5):

See let me help you for first one and you would be able to find out the answer to last one.

OpenStudy (vishweshshrimali5):

See: (1) f(x) = x/3 and g(x) = 1/x. Now, you are asked to find out f(x) - g(x). So, f(x) - g(x) can be calculated by just putting given value of f(x) in place of g(x) and value of g(x) in place of g(x). So, f(x) - g(x) = x/3 - 1/x. Did you get this ?

OpenStudy (anonymous):

yes i remember that portion..only to that part though other than that i get confused

OpenStudy (anonymous):

x